# Nearest Neighbor Analysis as a New Probe for Fuzzy Dark Matter

### v1

@@ Abstract
pagerange: Nearest Neighbor Analysis as a New Probe for Fuzzy Dark MatterA
@@ 1 Introduction
This paper is organized as follows: In Sec. 2, we review the basic concepts and formulas related to fuzzy dark matter and the nearest neighbor analysis. In this section, we introduce the functions and statistical parameters that we use in our analysis. Sec. 3, describes the various simulations we employ in this paper. Then, in Sec. 4, we utilize the nearest neighbor analysis to investigate the simulations and compute and compare for them the various functions and moments described in Sec. 2. Finally, in Sec. 5, after summarizing our key findings, we conclude by offering some additional remarks. Supplementary material, in particular, some plots are put in Appendix A.
@@ 2.2.1 SC-CDF/PDF & NN-CDF/PDF
Since the under-dense regions in the universe (voids) are significantly larger than overdense regions (e.g., filaments, halos, and others.), the random points chosen to calculate the SC-CDF (or SC-PDF) are most likely to be located in the voids, and the distance to their neighbors mainly depends on the size and the shape of the large-scale voids. So, the SC-CDF/PDF is a function that mainly probes the large scales and the general shape of the cosmic web. In contrast, the NN-CDF/PDF probes the small scales since the points and their nearest neighbors are chosen from the primary data points (e.g., halos), generally placed relatively close distances in the dense structures. So, it is more sensitive to the clustering of structures. As we will see shortly, all ${\mathbf{n}}$-point correlation functions influence SC-CDF/PDF and NN-CDF/PDF; as a result, they have information beyond the two-point correlation function. In this paper, the SC-CDF is denoted by $F(r)$ while the NN-CDF is shown by $G(r)$. We also use another appropriate function called $J$-function, defined as (Kerscher et al., 1999):
@@ 2.2.2 NNA
We introduce the probability distribution function of the angular distance between the first and second neighbors as a novel quantity to probe the cosmic web. This function is sensitive to the shape of the cosmic web and can be used to compare the predictions of different models of the LSS and DM (Ansarifard, 2024). As the first step, we must identify a data point’s first and second neighbors to estimate this function. Then, we calculate the angle $\phi$ between the two lines emanating from a data point to its nearest neighbors and call it NN-angle (NNA, see Fig. 1). Finally, we will repeat it for all data points to find the probability distribution function of angle separation of nearest neighbors. Notably, the NNA analysis probes the cosmic web’s shape, insensitive to conformal transformation or rescaling length scales.
@@
As matter clustering becomes stronger, the PDF of NNA near the two extremes where $\cos\phi\simeq\pm 1$ increases. This results in a larger $a_{4}$ in the fitting formula. As a result, during the NNA analysis, we will focus on the $a_{4}$ parameter in addition to the mean and standard deviation of the PDF of NNA with respect to $\phi$.
@@ 3 Cosmological Simulations
We use the Nearest Neighbor analysis, a simple yet powerful tool, to study cosmic structures under different theoretical models; specifically, we compare the above statistical parameters. For this purpose, we utilize data gathered from numerous FDM and CDM simulations. Additionally, we analyze whether the NN analysis can discriminate the influence of FDM’s QP and baryonic feedback on the cosmic web by comparing data from a simulation that incorporates baryonic effects and FDM simulations. For the sake of future comparison with observational data, it’s better to consider the positions of dark matter halos instead of simulation particles for NN analysis. So, we work with the halo catalogs of desired simulations. The NN-analysis’s results slightly depend on the mass ranges of the studied halos (see, Ansarifard et al., 2022). To ensure comparability among the results obtained from various simulations, we only consider halos within the mass range $M/M_{\odot}=\left[7\times 10^{7},7\times 10^{10}\right]$.
@@
    Figure 2: The probability distribution functions of spherical contact (or SC-PDF) for FDM, CDM, TNG, and TNG-Dark simulations, as a function of dimensionless parameter x (defined in Eq. 20), which denotes the normalized scale. As discussed in the text, due to the weakness of correlations in FDM relative to CDM, its SC-PDF in small x-scales is higher than others.
@@
    Figure 3: Probability distribution functions of nearest neighbor distances (or NN-PDF) for FDM, CDM, TNG, and TNG-Dark simulations, as a function of dimensionless parameter $x$. The suppression in small x-scales can be seen for FDM. This is related to suppressing the formation of smaller halos, generally located closer to other halos.
@@ 4 Analysis
As mentioned, this study involves seven simulations, including two CDM simulations, four FDM simulations, and one CDM simulation with baryonic effects. In this section, we will calculate and list the moments described in Sec 2.2 for all these simulations. To keep things simple, we will only present plots for one of the FDM simulations (simulation 4 in Table 1) and compare it with other simulations, including CDM, TNG50, and TNG50-Dark. The plots for the other three simulations can be found in the Appendix A. Notably, these plots align perfectly with the conclusions drawn from the simulation (4).
In the following, we first conduct the NN distance and angle analyses.
Figure 4: Functions $F(x)$, $G(x)$, and $J(x)$ against the dimensionless $x$ parameter (defined in Eq. (13)) for various simulations, with error bars displayed for each. The values of these functions are nearly identical for the CDM, TNG50, and TNG50-Dark simulations, but for the FDM simulation, they exhibit significant deviations from the values for the other simulations. This is due to the suppression in clustering and the subsequent weakness of correlation functions for the FDM simulation, which lead to higher values of $F(x)$ and $J(x)$ and a lower value of $G(x)$ relative to the other three simulations. At large scales, the error bars of $J(x)$ become very large as the values of $G(x)$ and $F(x)$ approach unity due to their particular functionality.
4.1 Nearest Neighbor’s Distance Analysis
[…]
Table 1: Moments of SC-PDF
@@ 4.1 Nearest Neighbor’s Distance Analysis
Table 2: Logarithmic moments of NN-PDF
@@
Tables 1-2 represent the values of the moments and the logarithmic moments, described in Sec.2.2, for our different simulations  [^7] The error bars are estimated by the Jacknife method (see, e.g., McIntosh, 2016). A noticeable difference is observed between the values obtained from the four FDM simulations and the other three simulations in all eight moments. The moments $s_{1}$, $s_{2}$, and $s_{3}$ show significant differences from each other ($\gtrsim 20$ percent), as do the logarithmic moments $l_{1}$, $l_{2}$, and $l_{3}$. On the other hand, $s_{4}$ and $l_{4}$ are slightly less sensitive to FDM short-scale effects.
@@ 4.2 Nearest Neighbors’ Angle Analysis
The NNA-PDF of our simulations differs from that of randomly distributed points due to the structure formation. Since the QP smears out the cosmic web on small scales, the PDF of NNA for FDM simulation is closest to that of a random distribution. Simulations of the CDM universe show no significant difference in the PDF of NNA with or without baryonic effects. Although the PDF of NNA for all simulations is more or less similar to a sine function, the maximum, mean, and other statistical quantities vary. Notably, the probability of very low and very high angles is higher in all simulations than in random points. During the formation of structures in the cosmic web, a large fraction of halos ($\sim 30-50$ percent [^8] There are discrepancies between studies as they use different criteria to segment the cosmic web (see, e.g., Cautun et al., 2014). See also Dome et al. (2023) for a comparison of the relative importance of cosmic nodes, filaments, walls and voids in CDM and FDM simulations.) reside in elongated, thin regions called filaments (Hahn et al., 2007; Cautun et al., 2014). This increases the chance of a halo and its first and second neighbors aligning in a nearly straight line. As shown in Fig. 5, the PDF of NNA for the FDM simulation is closer to that of random points. This can be attributed to the distinct structure of FDM filaments, which differs significantly from that of CDM simulations. In CDM simulations, filaments consist of many separate (sub-)halos, even on the smallest scales. However, FDM filaments do not break up into small halos. Instead, they form large, thick, and dense structures that connect large halos to each other. (see, e.g., Mocz et al., 2019, 2020; May & Springel, 2021, 2023). Accordingly, FDM filaments have fewer small halos, resulting in a lower probability of three neighboring halos being aligned in a straight line in FDM simulations.
Figure 5: The probability distribution function (PDF) for the angular distance of the first and second nearest neighbors (NNAs) of halos is shown for various simulations. The PDF of NNA for a set of random points is also depicted for comparison. It behaves like a re-scaled sine function. In the FDM simulation, the probability of this angle being nearly $90^{\circ}$ is higher than the other simulations, while the probabilities for the two opposite extremes ($0^{\circ}$ and $180^{\circ}$) are lower. Therefore, the PDF for FDM is the most similar to the sine function across different simulations. This is because the filaments in an FDM universe are blurred and contain fewer halos than a CDM universe.
@@
The simulations’ probability density function $P(\cos\phi)$ can be approximated well with quartic functions, as shown in Fig. 7. Table 3 presents the mean and standard deviation (st.d.) of NNA, along with the quartic coefficient $a_{4}$, which characterizes the convexity of $P(\cos\phi)$ due to clustering. The table also compares these parameters with a sample of randomly distributed points for reference.
Figure 6: The probability distribution function (PDF) for the cosine of the angle between the first and second nearest neighbors (NNAs) of halos is shown for various simulations. The PDF of NNA as a function of $\cos{\phi}$ for a set of random points is also depicted for comparison, which is nearly flat. The probability of $\cos{\phi}$ being zero is lower than the two extremes ($\cos{\phi}=\pm 1$). The apparent contradiction between this and what can be seen in Fig.5 is due to the fact that the bins near the two extremes correspond to a wider range of angles than the other bins. In a FDM universe, the filaments are blurred and contain fewer halos than in a CDM universe. This leads to a more uniform distribution of angles between the NNAs of halos, resulting in a flatter PDF.
In the case of a completely random spatial distribution of points, where no clustering is present, the probability density function (PDF) of nearest neighbor angular distances $P(\cos\phi)=1/2$ and $P(\phi)\propto\sin\phi$. As listed in Table 3, in agreement with analytic predictions, the average value of $\phi$ is $90^{\circ}$, the standard deviation is about $39^{\circ}$, and $a_{4}\simeq 0$.
Table 3: Nearest Neighbors’ Angle Analysis Moments
@@
The first three simulations produced significantly different results compared to the following four FDM simulations. In the FDM simulations, the NNAs are more closely grouped around the central angle of approximately 90 degrees. As a result, the mean angle is higher and closer to $90^{\circ}$, and the standard deviation is smaller, indicating that the NNAs are more similar to random points. This is because the distinct structure of FDM filaments leads to more randomly distributed halos and significantly lower $a_{4}$ values, indicating weaker clustering in FDM simulations.
The simulation of FDM that has the lowest mass has values of the standard deviation of $\phi$ and $a_{4}$ that are very similar to the values obtained for the sample of random points. This observation, when combined with the fact that the lower FDM mass leads to stronger suppression of the small-scale structures, indicates that these parameters are highly sensitive to the fuzzy nature of dark matter. On the contrary, as it is listed in Table 3, the mean of $\phi$ does not exhibit any difference for different FDM masses. This indicates that using the average of $\phi$ may not be effective in discriminating FDM and CDM physics.
Comparing the values of $a_{4}$ and the standard deviation of $\phi$ for simulation (7) with simulations (4) and (5) as listed in Table. 3, reveals that using FDM ICs provides closer values to those obtained for the random points. This highlights the effectiveness of these two parameters for investigating the fuzzy nature of dark matter. However, the $a_{4}$ parameter is more promising due to larger relative differences observed among the different simulations. Therefore, the NNA analysis captures information about the structures of the cosmic web in an FDM universe, proving a promising probe for FDM.
@@ 5 Summary and Conclusion
We consistently noticed a clear and distinct difference between the FDM simulations and the others during our analysis. The data in Tables 1-3 and Figs. 8-13 in Appendix A show that changing the resolution and using the FDM IC does not change this situation. Moreover, in most cases, reducing the FDM mass resulted in a greater deviation from the CDM simulations. The TNG simulation incorporates baryonic effects and its results were compared to other CDM simulations. Despite the difference in the two types of simulations, the NN analysis was able to accurately distinguish between their statistical properties. This indicates that the NN analysis can serve as a reliable tool for detecting FDM effects and differentiating between baryonic and fuzzy effects. These findings highlight the effectiveness of NN analysis in probing and identifying FDM effects.
@@
Furthermore, the chosen mass range for the halos ($\left[7\times 10^{7},7\times 10^{10}\right]M_{\odot}$) corresponds to galaxy mass scales for which there is limited observational data available.

### v2

@@ Abstract
pagerange: Nearest Neighbor Analysis as a New Probe for Fuzzy Dark Matter14
@@ 1 Introduction
This paper is organized as follows: In Sec. 2, we review the basic concepts and formulas related to fuzzy dark matter and the nearest neighbor analysis. In this section, we introduce the functions and statistical parameters that we use in our analysis. Sec. 3, describes the various simulations we employ in this paper. Then, in Sec. 4, we utilize the nearest neighbor analysis to investigate the simulations and compute and compare for them the various functions and moments described in Sec. 2. Finally, in Sec. 5, after summarizing our key findings, we conclude by offering some additional remarks. Supplementary material, in particular, some plots are put in Appendix B.
@@ 2.2.1 SC-CDF/PDF & NN-CDF/PDF
Since the under-dense regions in the universe (voids) are significantly larger than overdense regions (e.g., filaments, halos, and others.)(see, e.g., Dome et al., 2023), the random points chosen to calculate the SC-CDF (or SC-PDF) are most likely to be located in the voids, and the distance to their neighbors mainly depends on the size and the shape of the large-scale voids. So, the SC-CDF/PDF is a function that mainly probes the large scales and the general shape of the cosmic web. In contrast, the NN-CDF/PDF probes the small scales since the points and their nearest neighbors are chosen from the primary data points (e.g., halos), generally placed relatively close distances in the dense structures. So, it is more sensitive to the clustering of structures. As we will see shortly, all ${\mathbf{n}}$-point correlation functions influence SC-CDF/PDF and NN-CDF/PDF; as a result, they have information beyond the two-point correlation function. In this paper, the SC-CDF is denoted by $F(r)$ while the NN-CDF is shown by $G(r)$. We also use another appropriate function called $J$-function, defined as (Kerscher et al., 1999):
@@ 2.2.2 NNA
We introduce the probability distribution function of the angular distance between the first and second neighbors as a novel quantity to probe the cosmic web. This function is sensitive to the shape of the cosmic web and can be used to compare the predictions of different models of the LSS and DM (Ansarifard, 2024). As the first step, we must identify a data point’s first and second neighbors to estimate this function. Then, we calculate the angle $\phi$ between the two lines emanating from a data point to its nearest neighbors and call it NN-angle (NNA, see Fig. 1). Finally, we will repeat it for all data points to find the probability distribution function of angle separation of nearest neighbors. Notably, the NNA analysis probes the cosmic web’s shape, insensitive to conformal transformation or rescaling length scales. figureb
@@
As matter clustering becomes stronger, the PDF of NNA near the two extremes, where $\cos\phi\simeq\pm 1$, increases. This results in a larger $a_{4}$ in the fitting formula. As a result, during the NNA analysis, we will focus on the $a_{4}$ parameter in addition to the mean and standard deviation of the PDF of NNA with respect to $\phi$.
@@ 3 Cosmological Simulations
We use the Nearest Neighbor analysis, a simple yet powerful tool, to study cosmic structures under different theoretical models; specifically, we compare the above statistical parameters. For this purpose, we utilize data gathered from numerous FDM and CDM simulations. Additionally, we analyze whether the NN analysis can discriminate the influence of FDM’s QP and baryonic feedback on the cosmic web by comparing data from a simulation that incorporates baryonic effects and FDM simulations. For the sake of future comparison with observational data, it’s better to consider the positions of dark matter halos instead of simulation particles for NN analysis. So, we work with the halo catalogs of desired simulations. The NN-analysis’s results slightly depend on the mass ranges of the studied halos (see, Ansarifard et al., 2022). To ensure comparability among the results obtained from various simulations, we only consider halos within the mass range $M/M_{\odot}=\left[7\times 10^{7},1\times 10^{11}\right]M_{\odot}$.
@@ 4 Analysis
As mentioned, this study involves seven simulations, including two CDM simulations, four FDM simulations, and one CDM simulation with baryonic effects. All of the halo catalogs used in this analysis belong to the redshift $z=3$ (See, Appendix A for higher redshifts). In this section, we will calculate and list the moments described in Sec 2.2 for all these simulations. To keep things simple, we will only present plots for one of the FDM simulations (simulation 4 in Table 1) and compare it with other simulations, including CDM, TNG50, and TNG50-Dark. The plots for the other three simulations can be found in the Appendix B. Notably, these plots align perfectly with the conclusions drawn from the simulation (4).
Table 1: Moments of SC-PDF at $z=3$
@@
Figure 2: The probability distribution functions of spherical contact (or SC-PDF) for FDM, CDM, TNG, and TNG-Dark simulations at $z=3$, as a function of dimensionless parameter x (defined in Eq. 20), which denotes the normalized scale. As discussed in the text, due to the weakness of correlations in FDM relative to CDM, its SC-PDF in small x-scales is higher than others.
In the following, we first conduct the NN distance and angle analyses.
4.1 Nearest Neighbor’s Distance Analysis
Figure 3: Probability distribution functions of nearest neighbor distances (or NN-PDF) for FDM, CDM, TNG, and TNG-Dark simulations at $z=3$, as a function of dimensionless parameter $x$. The suppression in small x-scales can be seen for FDM. This is related to suppressing the formation of smaller halos, generally located closer to other halos.
Figure 4: Functions $F(x)$, $G(x)$, and $J(x)$ against the dimensionless $x$ parameter (defined in Eq. (13)) for various simulations at $z=3$, with error bars displayed for each. The values of these functions are nearly identical for the CDM, TNG50, and TNG50-Dark simulations, but for the FDM simulation, they exhibit significant deviations from the values for the other simulations. This is due to the suppression in clustering and the subsequent weakness of correlation functions for the FDM simulation, which lead to higher values of $F(x)$ and $J(x)$ and a lower value of $G(x)$ relative to the other three simulations. At large scales, the error bars of $J(x)$ become very large as the values of $G(x)$ and $F(x)$ approach unity due to their particular functionality.
[…]
Table 2: Logarithmic moments of NN-PDF at $z=3$
@@ 4.1 Nearest Neighbor’s Distance Analysis
Tables 1-2 represent the values of the moments and the logarithmic moments, described in Sec.2.2, for our different simulations  [^7] The error bars are estimated by the Jackknife method (see, e.g., McIntosh, 2016) throughout this paper. A noticeable difference is observed between the values obtained from the four FDM simulations and the other three simulations in all eight moments. The moments $s_{1}$, $s_{2}$, and $s_{3}$ show significant differences from each other ($\gtrsim 20$ percent), as do the logarithmic moments $l_{1}$, $l_{2}$, and $l_{3}$. On the other hand, $s_{4}$ and $l_{4}$ are slightly less sensitive to FDM short-scale effects.
@@ 4.2 Nearest Neighbors’ Angle Analysis
The NNA-PDF of our simulations differs from that of randomly distributed points due to the structure formation. Since the QP smears out the cosmic web on small scales, the PDF of NNA for FDM simulation is closest to that of a random distribution. Simulations of the CDM universe show no significant difference in the PDF of NNA with or without baryonic effects. Although the PDF of NNA for all simulations is more or less similar to a sine function, the maximum, mean, and other statistical quantities vary. Notably, the probability of very low and very high angles is higher in all simulations than in random points. During the formation of structures in the cosmic web, a large fraction of halos ($\sim 30-50$ percent [^8] There are discrepancies between studies as they use different criteria to segment the cosmic web (see, e.g., Cautun et al., 2014). See also Dome et al. (2023) for a comparison of the relative importance of cosmic nodes, filaments, walls and voids in CDM and FDM simulations.) reside in elongated, thin regions called filaments (Hahn et al., 2007; Cautun et al., 2014). This increases the chance of a halo and its first and second neighbors aligning in a nearly straight line. As shown in Fig. 5, the PDF of NNA for the FDM simulation is closer to that of random points. This can be attributed to the distinct structure of FDM filaments, which differs significantly from that of CDM simulations. In CDM simulations, filaments consist of many separate (sub-)halos, even on the smallest scales. However, FDM filaments do not break up into small halos. Instead, they form large, thick, and dense structures that connect large halos to each other. (see, e.g., Mocz et al., 2019, 2020; May & Springel, 2021, 2023). Accordingly, FDM filaments contain fewer halos than CDM filaments(Dome et al., 2023), resulting in a lower probability of three neighboring halos being aligned in a straight line in FDM simulations.
Figure 5: The probability distribution function (PDF) for the angular distance of the first and second nearest neighbors (NNAs) of halos is shown for various simulations at $z=3$. The PDF of NNA for a set of random points is also depicted for comparison. It behaves like a re-scaled sine function. In the FDM simulation, the probability of this angle being nearly $90^{\circ}$ is higher than the other simulations, while the probabilities for the two opposite extremes ($0^{\circ}$ and $180^{\circ}$) are lower. Therefore, the PDF for FDM is the most similar to the sine function across different simulations. This is because the filaments in an FDM universe are blurred and contain fewer halos than a CDM universe.
@@
Figure 6: The probability distribution function (PDF) for the cosine of the angle between the first and second nearest neighbors (NNAs) of halos is shown for various simulations at $z=3$. The PDF of NNA as a function of $\cos{\phi}$ for a set of random points is also depicted for comparison, which is nearly flat. The probability of $\cos{\phi}$ being zero is lower than the two extremes ($\cos{\phi}=\pm 1$). The apparent contradiction between this and what can be seen in Fig.5 is due to the fact that the bins near the two extremes correspond to a wider range of angles than the other bins. In a FDM universe, the filaments are blurred and contain fewer halos than in a CDM universe. This leads to a more uniform distribution of angles between the NNAs of halos, resulting in a flatter PDF.
@@
We see that a quartic function is a good fit to the probability density function $P(\cos\phi)$, as demonstrated in Fig. 7. However, due to the flattened shape of the distribution, different models can be distinguished by the fourth-order coefficient $a_{4}$. In other words, in our analysis, we see that the curvature of the distribution function is mainly controlled by the fourth-order coefficient, i.e., $a_{4}$ (see, Eq. 24). Therefore, $a_{4}$ proves to be a good measure of clustering.
Table 3: Nearest Neighbors’ Angle Analysis Moments at $z=3$
@@
Table 3 presents the mean and standard deviation (st.d.) of NNAs, along with the quartic coefficient $a_{4}$. The table also compares these parameters with a sample of randomly distributed points for reference. In the case of a completely random spatial distribution of points, where no clustering is present, we have $P(\cos\phi)=1/2$ and $P(\phi)\propto\sin\phi$. For these random points, as listed in Table 3, in agreement with simple analytic calculation for a completely random distribution, the average value of $\phi$ is $90^{\circ}$, the standard deviation is about $39^{\circ}$, and $a_{4}\simeq 0$.
The first three (CDM) simulations give significantly different results than the subsequent four(FDM) simulations. In the FDM simulations, the NNAs are more closely grouped around the central angle of about 90 degrees(See Fig. 6). As a result, the mean angle is higher and closer to $90^{\circ}$, and the standard deviation is less. Also, the $a_{4}$ parameter is lower for the FDM simulations, again indicating that the NNAs for the FDM simulations are more similar to random points. That is because the weaker clustering in the FDM simulations and the lack of halos in the filaments lead to more randomly distributed halos, resulting in significantly lower $a_{4}$. This observation supports the choice of $a_{4}$ as a measure of clustering.
The lower FDM mass leads to a larger de-Broglie wavelength and stronger suppression of the small-scale structures, resulting in fewer halos in filaments. Thus, we expect the NNA statistical parameters for the lower FDM masses to be more similar to the random points. The results in Table 3 are in full agreement with this expectation, as the lowest mass FDM simulation has values for the standard deviation of $\phi$ and $a_{4}$ that are closer to the values obtained for the random points sample compared to the other FDM simulations. This suggests that these two parameters are very sensitive to the fuzzy nature of dark matter.
As shown in Table 3, there is no noticeable variation in the average value of $\phi$ with respect to different FDM masses. This implies that using the mean of $\phi$ alone may not be a reliable and efficient parameter to differentiate between FDM and CDM physics.
Comparing the values of $a_{4}$ and the standard deviation of $\phi$ for simulation (7) with simulations (4) and (5) as listed in Table. 3, reveals that using FDM ICs provides closer values to those obtained for the random points. This, again, highlights the effectiveness of these two parameters for investigating the fuzzy nature of dark matter. However, the $a_{4}$ parameter is more promising due to larger relative differences observed among the different simulations. Therefore, the NNA analysis captures information about the structures of the cosmic web in an FDM universe, proving a promising probe for FDM.
@@ 5 Summary and Conclusion
We consistently noticed a clear and distinct difference between the FDM simulations and the others during our analysis. The appendix A shows that this is also true at higher redshifts, and becomes even more pronounced. The data in Tables 1-3 and Figs. 9-14 in Appendix B show that changing the resolution and using the FDM IC does not change this situation. Moreover, in most cases, reducing the FDM mass resulted in a greater deviation from the CDM simulations. The TNG simulation incorporates baryonic effects and its results were compared to other CDM simulations. We found no significant difference between them. So, the NN analysis is also able to distinguish between the statistical properties of FDM and baryonic simulations accurately. This indicates that the NN analysis can serve as a reliable tool for detecting FDM effects and differentiating between baryonic and fuzzy effects. These findings highlight the effectiveness of NN analysis in probing and identifying FDM effects.
@@
Furthermore, the chosen mass range for the halos ($\left[7\times 10^{7},1\times 10^{11}\right]M_{\odot}$) corresponds to galaxy mass scales for which there is limited observational data available.